OpenStudy (meilendurcer):

11/15 + 1/8 A. about 1/2 B. about 3/4 C. about 1 D. about 1 1/2

OpenStudy (anonymous):

make denominators common then add

OpenStudy (anonymous):

nphuongsun93 is right, Find the greatest common denominator.

OpenStudy (anonymous):

\[\frac{ 11 }{ 15 }+\frac{ 1 }{ 8 } = \frac{ (11\times8)+(15\times1) }{15\times8 }=...\] You can do the rest, right?

OpenStudy (meilendurcer):

@Husnul_Aini no but i thought i could.

OpenStudy (anonymous):

I'm sure you can do it ^^

OpenStudy (oaktree):

OK. So, you want to get common denominators. If you look at what Husnul_Aini wrote, then you can simply reduce the fraction.

OpenStudy (oaktree):

First, the top.

OpenStudy (oaktree):

11 x 8 = ?

OpenStudy (oaktree):

@meilendurcer?

OpenStudy (meilendurcer):

@OakTree 88

OpenStudy (oaktree):

Good. Then,15 x 1 is?

OpenStudy (meilendurcer):

15

OpenStudy (oaktree):

88+15=?

OpenStudy (meilendurcer):

103

OpenStudy (oaktree):

Yup. Then, the bottom is 15 x 8, or...?

OpenStudy (meilendurcer):

11x1?

OpenStudy (oaktree):

Where did you get that? What is 15 x 8?

OpenStudy (oaktree):

Just multiply.

OpenStudy (meilendurcer):

120

OpenStudy (oaktree):

There you go. So, your answer is 103/120.

OpenStudy (meilendurcer):

thanks